# A Comprehensive Survey on Data Augmentation

### v2

@@
Abstract.
Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples. By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities. Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process. To bridge this gap, we propose a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities. Specifically, from a data-centric perspective, this survey proposes a modality-independent taxonomy by investigating how to take advantage of the intrinsic relationship between data samples, including single-wise, pair-wise, and population-wise sample data augmentation methods. Additionally, we categorize data augmentation methods across five data modalities through a unified inductive approach.
@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
@@
With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition (He et al., 2016) and text translation (Bahdanau et al., 2015) to complicated scientific research tasks such as protein structure prediction (Jumper et al., 2021) and weather forecasting (Bi et al., 2023). Recently, powered by Latent Diffusion Models (Rombach et al., 2022) and Transformer (Vaswani et al., 2017), Stable Diffusion (Rombach et al., 2022) and ChatGPT (OpenAI, 2023) have greatly changed human beings’ working, living, and entertainment. The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations. The performances of AI models will be affected by the quantity and quality of training data, e.g. models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with unbalanced samples will lead to poor generalization ability. In most situations, researchers have to overcome difficulties caused by limited data sample amounts and imbalanced distribution of data samples. An intuitive solution to these problems is to acquire more data, but data are scarce or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
@@
To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient (Zha et al., 2023). The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data. It also introduces more diversity and fills the gap between training datasets and real-world applications. As technology continues to evolve, a variety of techniques are applied to augment data samples. Some simple approaches only randomly mask part of the data (DeVries and Taylor, 2017b; Zhong et al., 2020; Wei and Zou, 2019), while more sophisticated augmentation schemes involve generative adversarial networks (Bowles et al., 2018) or reinforcement learning agents (Cubuk et al., 2019).
@@
An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation. Existing literature surveys summarize data augmentation methods from different perspectives (Mikołajczyk and Grochowski, 2018; Shorten and Khoshgoftaar, 2019; Khalifa et al., 2022; Mumuni and Mumuni, 2022; Feng et al., 2021; Shorten et al., 2021; Bayer et al., 2022; Ding et al., 2022; Zhao et al., 2022; Zhou et al., 2022; Borisov et al., 2022; Sauber-Cole and Khoshgoftaar, 2022; Fonseca and Bacao, 2023; Wen et al., 2021; Iglesias et al., 2023; Iwana and Uchida, 2021). As summarized in Table. 1, each of them concentrates on a certain data modality. Most of them categorize data augmentation methods with modality-specific (e.g. image modality and text modality), or operation-centric (e.g. feature scaling operation, data perturbation operation) taxonomies. These surveys reflect the development of data augmentation and its application in different learning scenarios. However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation. To fill this gap, we investigate data augmentation methods comprehensively, across five most popular data modalities. We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information. Then, we propose a unified taxonomy that focuses on the way data augmentation derives new data samples. Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
The main contributions of this survey can be summarized as follows:
@@
  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation.
@@
Table 1. A comparison of data modalities and taxonomies discussed by existing surveys and ours.
Data Modality Taxonomies
Image Text Graph Tabular Time-series Modality Technical Procedural Sample amount Information type
IIPhDW 2018 (Mikołajczyk and Grochowski, 2018)
J Big Data 2019 (Shorten and Khoshgoftaar, 2019)
Artif Intell Rev 2022 (Khalifa et al., 2022)
Array 2022 (Mumuni and Mumuni, 2022)
IJCNLP 2021 (Feng et al., 2021)
J Big Data 2021 (Shorten et al., 2021)
Comput Surv 2022 (Bayer et al., 2022)
KDD 2022 (Ding et al., 2022)
arXiv 2022 (Zhao et al., 2022)
arXiv 2022 (Zhou et al., 2022)
TNNLS 2022 (Borisov et al., 2022)
J Big Data 2022 (Sauber-Cole and Khoshgoftaar, 2022)
J Big Data 2023 (Fonseca and Bacao, 2023)
IJCAI 2021 (Wen et al., 2021)
Neural Comput Appl 2023 (Iglesias et al., 2023)
Plos one 2021 (Iwana and Uchida, 2021)
@@
2. Background
2.1. Early Evolution
@@
Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early. The concept of data augmentation was conceived at the very beginning of Deep Learning. For example, the use of random distortion can be found in LeNet (LeCun et al., 1998), where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error. Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) (Simard et al., 2003; Cireşan et al., 2011; Ciregan et al., 2012). AlexNet (Krizhevsky et al., 2012) explicitly employs several data augmentation techniques to reduce overfitting. It augments the datasets by extracting patches from images and altering the color intensity. It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption. Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods (DeVries and Taylor, 2017b). These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it. SMOTE (Chawla et al., 2002) is another early and representative work on data augmentation but from a different perspective. Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
2.2. Existing Surveys
@@
Given the rapid development of CNNs and their wide application in image-related tasks, most early data augmentation methods are proposed for image data. Also, it is natural to augment image data because many data augmentation methods suit the traits of CNNs very well, such as translation invariance. Published in 2018, (Mikołajczyk and Grochowski, 2018) is one of the first surveys on data augmentation. It gives a brief discussion of the effect of traditional and neural-based augmentation methods. In 2019, the use of data augmentation on image data is thoroughly studied in (Shorten and Khoshgoftaar, 2019), which is up to now the most extensive work on this track. It categorizes image data augmentation methods into basic image manipulation and deep learning approaches, with a wide coverage. It also introduces how meta-learning is used to wisely choose and combine various basic operations for data augmentation. (Khalifa et al., 2022) is a recent survey on image data augmentation published in 2022. It formalizes commonly used operations for image data augmentation with equations and illustrates their effects with examples. In the same year, (Mumuni and Mumuni, 2022) also elaborates on image data augmentation with plenty of examples. It adopts taxonomies similar to those of (Shorten and Khoshgoftaar, 2019) but includes more modern approaches and evaluates the improvements in model performance when these augmentation methods are used.
@@
Data augmentation on text data is not thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components. Still, several augmentation methods are proposed for or transplanted to text data. (Feng et al., 2021) is among the first surveys on text data augmentation. It adopts a simple taxonomy but gives more emphasis on the applications. Built on previous work on image data augmentation, (Shorten et al., 2021) continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation. (Bayer et al., 2022) categorizes text data augmentation methods by identifying whether they are used on the data space or feature space. It also presents the performance improvements of different methods.
@@
Following the success of data augmentation methods for image and text data, there is a growing trend in augmenting graph data, too. Since the graph represents data in a relatively complex form compared with image and text, there are more opportunities for data augmentation. This also leads to more possible ways to categorize graph data augmentation methods. For example, (Ding et al., 2022) views graph data augmentation from the employed techniques and application scenarios, while (Zhao et al., 2022) discusses graph data augmentation in terms of its data modality, task level, and whether the augmentation operation is rule-based or learned. (Zhou et al., 2022) tries to elaborate on all feasible taxonomies for graph data augmentation and uses detailed schematic illustrations to explain the overview framework as well as the idea of some typical methods.
@@
Tabular data is common in reality, but there is not as much room for data augmentation as other data types because tabular data lacks some features for augmentation to utilize. Some surveys only discuss augmentation techniques in the context of deep learning on tabular data (Borisov et al., 2022), and others mainly focus on how to generate tabular data (Sauber-Cole and Khoshgoftaar, 2022; Fonseca and Bacao, 2023).
@@
With its growing application in scenarios such as speech recognition and IoT, time-series data is attracting more attention. (Wen et al., 2021) discusses basic and advanced augmentation methods. (Iglesias et al., 2023) focuses on generative approaches. (Iwana and Uchida, 2021) gives more discussion on mixing time-series data.
@@
Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data. Such a survey can be profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples. Summarizing these similarities can help reveal the common pattern of data augmentation. A comprehensive survey does mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other. Some taxonomies contrast basic operations to advanced approaches; some categorize methods by their application scenarios. Besides, these taxonomies provide limited insight into what happens to the information. Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or application-based taxonomies reflect the development of data augmentation techniques. However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation. Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and information-centric taxonomy.
3. Taxonomy
@@
(1) $f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$
This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, in this section, we only use a labeled dataset to explain the taxonomy. Some data augmentation approaches are used in the input data space, while others are used in the latent feature space. This section uses $\mathbf{X}$ as a space-agnostic representation.
In this comprehensive survey, we try to propose a two-tier taxonomy from an information-centric perspective that can be applied to all data modalities. The main consideration of our taxonomy is whence the information in the augmented data comes. It differentiates data augmentation techniques by asking two research questions: RQ1: How many samples are used to generate each new sample? RQ2: Which part of the information is used to generate new data? The answers to these questions constitute the taxonomy hierarchy which we present below.
By answering RQ1, we can divide data augmentation approaches into Individual, Multiple, and Populational Augmentation, which build the first tier of the taxonomy in our survey.
Individual Augmentation. This type of data augmentation method transforms one data sample without referring to others to create new data. It introduces a certain perturbation to the data, such as masking and noise addition. The information of new data comes from exactly one sample in the original dataset, and each new sample often lies around its original sample in the feature space. We formalize Individual Augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation that is related to the original data $\mathbf{x}_{i}$:
(2) $\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$
We further answer RQ2 to build the second tier of the taxonomy. To answer this question, we must consider how the information is contained in the data. Typically, data is made up of a certain type of element that carries unit information with a value, such as the coloring of a pixel or the wording in a sentence. For most data, there is also some structural information that connects these elements in a certain way, such as the positional relationship of pixels or the syntax of a sentence. The key to designing data augmentation methods is to think of how to properly leverage the two types of information. So, we can categorize Individual Augmentation according to which part of the information is used for data augmentation:
@@
Multiple Augmentation. This type of data augmentation method makes use of multiple data to acquire new ones. Information from different data samples is combined by operations such as interpolation and concatenation. New samples derived from these methods often lie between the source data in the feature space. We formalize Multiple Augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
(3) $\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$
Now we answer RQ2 for Multiple Augmentation. Similarly to Individual Augmentation, this type of data augmentation is also conducted utilizing either value of structure information, and the detailed classification is as follows:
@@
Populational Augmentation. This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the population of the data. It comprehensively learns the features and abstractions of the entire dataset and samples completely new data from a learned distribution of the data. The generated data should also fall into this distribution. We formalize Populational Augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
(4) $\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$
The answer to RQ2 for Populational Augmentation is different from the two categories mentioned above. When generating new data based on population distributions, most methods simultaneously consider value and structural information. Some methods also introduce external knowledge when learning from the data population. So, we have the following classification:
@@
  • Exogenous Generation, where the generation process refers to external resources such as other datasets or expert knowledge.
@@
We have presented a general taxonomy for data augmentation, whereas the category is at a high level, and each data modality has its specific data augmentation techniques depending on the nature of that modality. In addition, the equations that we use as a general formalization are only abstract and conceptual. They do not literally represent all specific data augmentation techniques in these categories. From paragraph 4 to paragraph 8, we will elaborate on how this taxonomy applies to each data modality and review some typical data augmentation methods.
Figure 1. A taxonomy for data augmentation (left) and how information in each data modality is utilized (right).
4. Data Augmentation for Image Data
The elemental components of an image are pixels. A pixel carries unit information that uses a numerical value to represent the color. Each pixel also has spatial relationships with other pixels, so, images are well-structured. Data augmentation for image data revolves around the manipulation of the colors and spatial relationships of image pixels.
4.1. Individual Image Augmentation
4.1.1. Value-based Image Transformation

### v3

@@
Abstract
Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples. By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities. Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process. To bridge this gap, this survey proposes a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities by investigating how to take advantage of the intrinsic relationship between and within instances. Additionally, it categorizes data augmentation methods across five data modalities through a unified inductive approach.
Index Terms:
@@ Abstract
I Introduction
@@
With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition [1] and text translation [2] to complicated scientific research tasks such as drug-drug interaction prediction [3] and single-cell RNA sequencing analysis [4]. Recently, powered by Latent Diffusion Models [5] and Transformer [6], Stable Diffusion [5] and ChatGPT [7] have greatly changed human beings’ working, living, and entertainment. The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations. The performances of AI models will be affected by the quantity and quality of training data, e.g. models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability. In most situations, researchers have to overcome difficulties caused by data scarcity and distribution imbalance. An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
@@
To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient [8]. The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data. It also introduces more diversity and fills the gap between training datasets and real-world applications. As technology continues to evolve, a variety of techniques are applied to augment data samples. Some simple approaches only randomly mask part of the data [9, 10, 11], while more sophisticated augmentation schemes involve generative adversarial networks [12] or reinforcement learning agents [13].
@@
An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation. Existing literature surveys summarize data augmentation methods from different perspectives [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. As summarized in Table. I, each of them concentrates on a certain data modality. Most of them categorize data augmentation methods with modality-specific (e.g. image modality and text modality) or operation-centric (e.g. feature scaling operation, data perturbation operation) taxonomies. These surveys reflect the development of data augmentation and its application in different learning scenarios. However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation. To fill this gap, we investigate data augmentation methods comprehensively, across the five most popular data modalities. We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information. Then, we propose a unified taxonomy that focuses on the way data augmentation leverages information from samples. Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
The main contributions of this survey are as follows:
@@
  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation;
@@
TABLE I: A comparison of data modalities and taxonomies discussed by existing surveys and ours.
Data Coverage Taxonomies
 Image Text Graph Tabular Time-series  Modality-specific Modality-independent Operation-centric Data-centric
IIPhDW 2018 [14]
J Big Data 2019 [15]
IJCNLP 2021 [18]
J Big Data 2021 [19]
IJCAI 2021 [27]
Plos one 2021 [29]
Artif Intell Rev 2022 [16]
Array 2022 [17]
Comput Surv 2022 [20]
KDD 2022 [21]
arXiv 2022 [22]
arXiv 2022 [23]
TNNLS 2022 [24]
J Big Data 2022 [25]
J Big Data 2023 [26]
Neural Comput Appl 2023 [28]
arXiv 2024 [30]
arXiv 2024 [31]
arXiv 2025 [32]
@@
II Background
II-A Early Evolution
@@
Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early. The concept of data augmentation was conceived at the very beginning of Deep Learning. For example, the use of random distortion can be found in LeNet [33], where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error. Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [34, 35]. AlexNet [36] explicitly employs several data augmentation techniques to reduce overfitting. It augments the datasets by extracting patches from images and altering the color intensity. It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption. Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods [9]. These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it. SMOTE [37] is another early and representative work on data augmentation but from a different perspective. Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
II-B Existing Surveys
@@
Given the rapid development of CNNs and their wide application in image-related tasks, most early data augmentation methods are proposed for image data. Also, it is natural to augment image data because many data augmentation methods suit the traits of CNNs very well, such as translation invariance. Published in 2018, [14] is one of the first surveys on data augmentation. It briefly discusses the effects of traditional and neural-based augmentation methods. In 2019, data augmentation usage on image data is thoroughly studied in [15]. It categorizes image data augmentation methods into basic image manipulation and deep learning approaches, with a wide coverage. It also introduces how meta-learning selects and combines basic data augmentation operations. [16] is a recent survey on image data augmentation published in 2022. It formalizes commonly used operations for image data augmentation with equations and illustrates their effects with examples. In the same year, [17] also elaborates on image data augmentation with plenty of examples. It adopts taxonomies similar to [15], but includes more modern approaches and evaluates the improvements in model performance.
@@
Data augmentation on text data has not been thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components. Still, several augmentation methods are proposed for or transplanted to text data. [18] is among the first surveys on text data augmentation. It adopts a simple taxonomy but gives more emphasis on the applications. Built on previous work on image data augmentation, [19] continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation. [20] categorizes text data augmentation methods by identifying whether they are used on the data space or feature space. It also presents the performance improvements of different methods. [31] focuses on the application of data augmentation for training Large Language Models. [32] discusses how Multimodal Large Language Models can be used for augmenting text, image, and speech data.
@@
Following the success of data augmentation methods for image and text data, there is a growing trend in augmenting graph data, too. Since the graph represents data in a relatively complex form compared with image and text, there are more opportunities for data augmentation. This also leads to more possible ways to categorize graph data augmentation methods. For example, [21] views graph data augmentation from the employed techniques and application scenarios, while [22] discusses graph data augmentation in terms of its data modality, task level, and whether the augmentation operation is rule-based or learned. [23] tries to elaborate on all feasible taxonomies for graph data augmentation and uses detailed schematic illustrations to explain the overview framework as well as the idea of some typical methods.
@@
Tabular data is common in reality, but there is not as much room for data augmentation as other data types because tabular data lacks some features for augmentation to utilize. Some early surveys only discuss augmentation techniques in the context of deep learning on tabular data [24], and others mainly focus on how to generate tabular data [25, 26]. Recently, [30] presented a dedicated and thorough review on tabular data augmentation, focusing on how generative models and retrieval techniques contribute to tabular data augmentation and discussing their future trends.
@@
With its growing application in scenarios such as speech recognition and IoT, time-series data is attracting more attention. [27] discusses basic and advanced augmentation methods. [28] focuses on generative approaches. [29] gives more discussion on mixing time-series data.
@@
Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data. Such a survey can be profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples. Summarizing these similarities can help reveal the common pattern of data augmentation. A comprehensive survey does not mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other. Some taxonomies contrast basic operations to advanced approaches; some categorize methods with modality-specific terms. Besides, these taxonomies provide limited insight into what happens to the information. Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or modality-specific taxonomies reflect the development of data augmentation techniques. However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation. Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and data-centric taxonomy.
III Taxonomy
@@
$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ (1)
This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}_{P}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, we only use a labeled dataset to explain the taxonomy. Some data augmentation approaches are used in the input data space, while others are used in the latent feature space. This section uses $\mathbf{X}$ as a space-agnostic representation.
In this survey, we try to propose a two-tier taxonomy from a data-centric perspective that can be applied to all data modalities. The main consideration of our taxonomy is whence the information in the augmented data comes. It differentiates data augmentation techniques by asking two research questions: RQ1: How many samples are used to generate each new sample? RQ2: Which part of the information is used to generate new data? The answers to these questions constitute the taxonomy hierarchy which we present below.
By answering RQ1, we can divide data augmentation approaches into Single-instance Level Augmentation, Multi-instance Level Augmentation, and Dataset Level Augmentation, which build the first tier of the taxonomy in our survey.
Figure 1: A taxonomy for data augmentation (left) and how information in each data modality is utilized (right).
Single-instance Level Augmentation. This type of data augmentation method transforms one data sample without referring to others to create new data. It introduces a certain perturbation to the data, such as masking and noise addition. The information of new data comes from exactly one sample in the original dataset, and each new sample often lies around its original sample in the feature space. We formalize single-instance level augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation related to the original data $\mathbf{x}_{i}$:
$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$ (2)
We further answer RQ2 to build the second tier of the taxonomy. To answer this question, we must consider how the information is contained in the data. Typically, data is made up of a certain type of element that carries unit information with a value, such as the coloring of a pixel or the wording in a sentence. For most data, there is also some structural information that connects these elements in a certain way, such as the positional relationship of pixels or the syntax of a sentence. The key to designing data augmentation methods is to think of how to wisely leverage the two types of information. So, we can categorize single-instance level augmentation according to which part of the information is used for data augmentation:
@@
Multi-instance Level Augmentation. This type of data augmentation method makes use of multiple data to acquire new ones. Information from different data samples is combined by operations such as interpolation and concatenation. New samples derived from these methods often lie between the source data in the feature space. We formalize multi-instance level augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$ (3)
Now we answer RQ2 for multi-instance level augmentation. Similarly to single-instance level augmentation, this type of data augmentation is also conducted utilizing either value or structure information, and the detailed classification is:
@@
Dataset Level Augmentation. This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the entire dataset. It comprehensively learns the features and abstractions of the dataset, and samples completely new data from a learned distribution of the dataset. The generated data should also fall into this distribution. We formalize dataset level augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
$\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$ (4)
The answer to RQ2 for dataset level augmentation is different from the two categories mentioned above. When generating new data based on dataset distributions, most methods simultaneously consider value and structural information. Some methods also introduce external knowledge when learning from the source dataset. So, we have the following classification:
@@
  • Exogenous Generation, which uses external resources like other datasets or expert knowledge.
@@
We have presented a general taxonomy for data augmentation at a high level with abstract and conceptual formalization. We organize this survey in accordance with this taxonomy, from Section IV to Section VI. We elaborate on how the proposed taxonomy applies to each data modality and review some typical data augmentation methods.
IV Single-instance Level Augmentation
Single-instance level augmentation methods perturb one data sample’s value or structure information to obtain the augmented data. Different data modalities represent value information with different elemental components and organize these components with different types of structure.
For image data, the elemental components are pixels. A pixel carries unit information that uses a numerical value to represent the color. Each pixel also has spatial relationships with other pixels, so, images are well-structured. Single-instance level augmentation for image data revolves around the perturbation of the image pixels’ colors and spatial relationships. For text data, the elemental components of text data are words (or sometimes, tokens). A word often expresses minimal meanings. Words are organized into sentences following syntactic rules. The sentence is a sequential representation since one word comes after another, but there is also a tree-structured grammatical dependency behind the sentence. Single-instance level augmentation for text data often changes the wording or reshapes the syntax of a sentence. A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ stands for the vertex set, and $\mathcal{E}$ stands for the edge set. The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}$ is $1$ if there is an edge connecting vertex $i$ and vertex $j$, and is 00 otherwise. In a deep graph learning scenario, the graph also contains a feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ that represents the attributes of the nodes. Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$. Single-instance level augmentation for graph data is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$. A table is made up of cells, which store numerical or categorical values that represent attributes of the described entities. Cells are organized into columns and rows, but this structural information is trivial, and switching the sequence of the columns and rows does not affect the information conveyed by the table. So, most data augmentation methods leverage the value information. Still, it is also possible to perturb the structure information of a table. Time-series data represent sequential values within a period, and each timestamp is considered an elemental component. Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.
That is, single-instance level augmentation methods focus on utilizing the values and structures of these components.
IV-A Value-based Transformation
IV-A1 Color Perturbation